Find the characteristic polynomial of the linear operator T:R^2 tends to R^2 defined by T(a, b)= (3a+5b, 2a-7b)
1
Expert's answer
2020-12-21T16:32:48-0500

Find the matrix A that represents T relative to the usual basis of "R^2" .we have

A= "\begin{bmatrix}
 3 & 5 \\
 2 & -7
\end{bmatrix}"

So we look for the trace of A.

tra(A)= 3+(-7) = -4

det(A) = -31

The Characteristics polynomial is.

"t^2-tr(A)t+|A|\\
\implies\\
t^2-(-4)t+(-31)\\
t^2+4t-31\\"